OpenStudy (anonymous):

solve for y in the first equation. x – 3y = –6 2x – 7y = 10

OpenStudy (anonymous):

its one of these answers ; 1/3x+2 - 1/3x+6 -x+2 -1/3x +2

OpenStudy (anonymous):

please help:) i don't want the answer i just want help

OpenStudy (anonymous):

write the given equation in the form of y=mx+c

OpenStudy (anonymous):

first, take x onto the right hand side.

jhonyy9 (jhonyy9):

so i think that you need to calcule the x from second equation and this resulted value of x you need just substitute in place of x in the first equation and in this way will get the y from the first equation

OpenStudy (anonymous):

let me try

jhonyy9 (jhonyy9):

hope so much that is understandably

jhonyy9 (jhonyy9):

so if you need to calcule the x from second equation what will be the first step ?

jhonyy9 (jhonyy9):

assum to both sides 7y

jhonyy9 (jhonyy9):

so than what will get ?

OpenStudy (anonymous):

don't i need to get y by itself? wouldn't it be - 2 on both sides?

jhonyy9 (jhonyy9):

because your exercise is to solve for y from the first equation ,yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than you need to get a value of x from the second equation and this value of x will need to substitute in place of x in the first equation ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than calcule please the first step how i have wrote it before little up what will get ?

OpenStudy (anonymous):

2x-7y=10 subtract 7y on both sides and get 2x=3 divide by 2 on each side and get 1.5

jhonyy9 (jhonyy9):

why ? how can you subtracting from 10 the 7y ?

OpenStudy (anonymous):

oh my word i'm confused, i got - 1/3x+6 as the answer is that correct?

jhonyy9 (jhonyy9):

please write it here step by step

jhonyy9 (jhonyy9):

adding to both sides 7y will get 2x=10+7y yes ?

jhonyy9 (jhonyy9):

so than from this how you will get the x ?

jhonyy9 (jhonyy9):

so do you like to continue it ?